# Beyond Matryoshka: Revisiting Sparse Coding for Adaptive Representation

### v2

@@ Abstract
Many large-scale systems rely on high-quality deep representations (embeddings) to facilitate tasks like retrieval, search, and generative modeling. Matryoshka Representation Learning (MRL) recently emerged as a solution for adaptive embedding lengths, but it requires full model retraining and suffers from noticeable performance degradations at short lengths. In this paper, we show that sparse coding offers a compelling alternative for achieving adaptive representation with minimal overhead and higher fidelity. We propose Contrastive Sparse Representation (CSR), a method that sparsifies pre-trained embeddings into a high-dimensional but selectively activated feature space. By leveraging lightweight autoencoding and task-aware contrastive objectives, CSR preserves semantic quality while allowing flexible, cost-effective inference at different sparsity levels. Extensive experiments on image, text, and multimodal benchmarks demonstrate that CSR consistently outperforms MRL in terms of both accuracy and retrieval speed-often by large margins-while also cutting training time to a fraction of that required by MRL. Our results establish sparse coding as a powerful paradigm for adaptive representation learning in real-world applications where efficiency and fidelity are both paramount. Code is available at this https URL.
@@
Figure 1: Overview of our proposed method. (a) Illustrative comparison between standard embeddings (dense, long) and two different compression schemes: Matryoshka representations (MRL) (Kusupati et al., 2022) with short length and our Contrastive Sparse Representation (CSR) based on sparsification. (b) Comparison of retrieval accuracy and time of different methods on ImageNet with GPU. Compared to MRL and int8 quantification (Quant Int8) methods, our sparse embedding approach CSR attains the best retrieval accuracy (very close to full representations) while being much more efficient in retrieval time, using sparse matrix multiplication on GPU. (c) Training GPU hours of CSR compared to baseline methods, where we outperform MRL on 1-NN accuracy with much less training time.
@@ 1 Introduction
Recently, a class of methods called Matryoshka Representation Learning (MRL) (Kusupati et al., 2022) has drawn a lot of attention and is now officially supported in the latest OpenAI and Google’s Gemini text embedding APIs  (OpenAI, 2024; Lee et al., 2024b) with millions of users and applications. The idea of MRL is to train an ensemble of representations truncated at different lengths (e.g., from 8 to 2048) through joint multi-task training. However, MRL deviates from standard representation learning and requires full parameter updates to the backbone; the joint training also inevitably sacrifices the quality of representations at a noticeable margin (e.g., 5% drop of top-1 accuracy on ImageNet at full representation length). These limitations render MRL a costly and lossy method for adaptive representation.
@@
Another key advantage of sparse features is that they eliminate the need to update the entire network. In contrast, MRL—Kusupati et al. (2022) noted—performs poorly unless full-parameter tuning. However, many existing foundation models, such as the multimodal representations in CLIP (Radford et al., 2021) and the text embeddings in NV-Embed (Lee et al., 2024a), are pre-trained as single representations on massive Internet-scale data. Finetuning these models would be relatively expensive compared to sparse methods. Leveraging recent advances in training sparse autoencoders (SAEs) (Cunningham et al., 2023; Gao et al., 2024), we can train a light-weight 2-layer MLP module for sparsifying pre-trained embeddings within a very short period of time (e.g., half of an hour on ImageNet with a single GPU), which is of orders of magnitude faster than MRL, as shown in Figure 1(c).
These pieces of evidence on accuracy, retrieval time, and training time show that sparse features are strong alternatives to MRL methods for producing high-fidelity and computationally efficient representations with a lightweight module and training cost. Our proposed method, Contrastive Sparse Representation Learning (CSR), combines contrastive retrieval and reconstructive autoencoding objectives to preserve the original feature semantics while better tailing it down to the retrieval tasks. We evaluate CSR on a range of standard embedding benchmarks, from image embedding, text embedding, to multimodal embeddings, and compare it against various state-of-the-art efficient embedding models. Extensive experiments show that CSR consistently outperforms MRL and its variants by significant margins in terms of both accuracy and efficiency. Notably, under the same compute budget, CSR rivals MRL’s performance by 17%, 15%, and 7% on ImageNet classification, MTEB text retrieval, and MS COCO retrieval, respectively. Our main contributions are:
@@ 2 Related Work
Recent research has increasingly focused on learning adaptive representations that cater to multiple downstream tasks with diverse computational requirements. Early efforts explored context-based architectural adaptations (Kim & Cho, 2020), dynamic widths and depths in BERT (Hou et al., 2020), and random layer dropping during training to improve pruning robustness (Fan et al., 2019). More recently, Matryoshka Representation Learning (Kusupati et al., 2022) introduced a novel technique for creating flexible, nested substructures within embeddings, enabling fine-grained control over the trade-off between latency and accuracy. This concept has since been extended to various modalities and applications, including large language models (OpenAI, 2024; Nussbaum et al., 2024; Yu et al., 2024), diffusion models (Gu et al., 2023), recommendation systems (Lai et al., 2024), and multimodal models (Cai et al., 2024; Hu et al., 2024). Other works have further explored token reduction in image and video processing (Yan et al., 2024b; Duggal et al., 2024).
Despite these advances, existing methods often do not fully harness the capabilities of large foundation models, highlighting the need for more effective compression strategies. Our proposed sparse coding methodology addresses this gap by providing a lightweight, plug-and-play solution that can be readily applied on top of any foundation model – significantly reducing computational overhead while preserving representational quality.
@@
Sparse coding serves as a powerful technique for compressing high-dimensional signals and extracting salient features (Wright et al., 2010; Zhang et al., 2015), with learned sparse representations often providing additional computational benefits. Prior work has induced sparsity through modifications to model architectures or training protocols, including modifications to attention mechanisms (Correia et al., 2019), applying $\ell_{1}$-norm penalties to neuron activations (Georgiadis, 2019), and promoting sparse activations in large language models (Mirzadeh et al., 2023; Zhang et al., 2024). However, training state-of-the-art foundation models from scratch under these sparsity constraints has proven challenging (Elhage et al., 2022), limiting their current applicability.
@@ 3 Method
Our proposed framework, Contrastive Sparse Representation learning (CSR), is illustrated in Figure 2. Starting from a pre-trained embedding $v\in\mathbb{R}^{d}$, we project it into a sparse representation space $\mathbb{R}^{h}$, selectively activating the most relevant dimensions for adaptive representation learning. We then regularize this hidden space using a reconstruction-based sparse compression loss (Section 3.2.1). Additionally, with theoretical motivations and guarantees provided by (Wang et al., 2024), we introduce a non-negative contrastive loss to expand model capacity.(Section 3.2.2)
@@ 3.1 Preliminaries
Matryoshka Representation Learning.
MRL simultaneously optimizes embeddings at multiple dimensions, as illustrated in Figure 2, to produce representations of variable size. Specifically, let $\mathcal{M}$ be a set of target embedding sizes. For each $m\in\mathcal{M}$, MRL applies an additional linear classifier to the first $m$ dimensions of the embedding vector, $v_{1:m}\in\mathbb{R}^{m}$. This design ensures each truncated representation is explicitly trained via the final loss. Formally, the MRL objective is:
$\mathcal{L}_{\mathrm{MRL}}=\sum_{m\in\mathcal{M}}c_{m}\mathcal{L}_{\mathrm{CE}}\left({\bm{W}}^{(m)}\cdot f(x_{i};\theta_{f})_{1:m};y_{i}\right)$ (1)
@@
3.2 Contrastive Sparse Representation Learning
As discussed in Section 1, Equation 1 faces two key constraints: it requires class labels $y_{i}$ and necessitates training the backbone parameters $\theta_{f}$. These limitations reduce its utility in scenarios where web-scale datasets often lack labels and fine-tuning large foundation models is prohibitively expensive. To overcome these constraints while preserving computational efficiency, we employ reconstruction-based sparse coding on top of pre-trained models and incorporate a task-specific loss for generalization ability.
3.2.1 Reconstruction-based sparse compression
Reconstruction-based sparse compression reduces data size by preserving only the most essential components, thereby enabling more efficient storage and computation. Specifically, we adopt sparse autoencoders (SAEs) due to their demonstrated ability to capture semantically rich features from foundation models (Cunningham et al., 2023; Yan et al., 2024a).
Sparse AutoEncoder.
SAEs aim to learn efficient data representations by reconstructing inputs while enforcing sparsity in the latent space. For an input embedding, vector $v:=f(x)\in\mathbb{R}^{d}$, we omit $\theta_{f}$ since SAEs do not optimize the backbone parameters. The encoder and decoder of SAE are defined by:
$z_{k}$ $:=\mathrm{TopK}({\bm{W}}_{\mathrm{enc}}(f(x)-\bm{b}_{\mathrm{pre}})+\bm{b}_{\mathrm{enc}})$ (2)
$\widehat{f(x)}_{k}$ $:={\bm{W}}_{\mathrm{dec}}z_{k}+\bm{b}_{\mathrm{pre}},$
[…]
$\mathcal{L}(k)$ $=\left\|f(x)-\widehat{f(x)}_{k}\right\|_{2}^{2}$ (3)
[…]
$\mathcal{L}_{\mathrm{recon}}=\mathcal{L}(k)+\mathcal{L}(4k)/8+\beta\mathcal{L}_{\mathrm{aux}}$ (4)
@@ 3.2.1 Reconstruction-based sparse compression
where $\mathcal{L}_{\mathrm{aux}}=||e-\hat{e}||^{2}_{2}$, $e=f(x)-\widehat{f(x)}$, and $\hat{e}=W_{\mathrm{dec}}z$ is the reconstruction using the top-$k_{\mathrm{aux}}$ dead latents. By default, we set $k_{\mathrm{aux}}=512$ and $\beta=1/32$, following the setting in (Gao et al., 2024). We also offer dynamic sparsity selection, with $k$ ranging from 8 to 256, to accommodate different tasks across various modalities.
[…]
3.2.2 Contrastive-based sparse compression
Furthermore, we combine task-specific losses to enhance adaptation to diverse downstream tasks. Specifically, we employ contrastive learning strategies for two key reasons: i) Unified loss function-it can integrate supervised tasks by treating intra-class data as positive samples and outer-class data as negative samples, following (Huang et al., 2024), while also being adaptable to unsupervised tasks such as retrieval. ii) Performance improvement-can be achieved by enhancing embedding capabilities through in-batch negative sampling with InfoNCE loss, as demonstrated by  (Carlsson et al., 2021; Zhang et al., 2020; Duan et al., 2024). The loss objective can be formulated as:
$\mathcal{L}_{\mathrm{cl}}=-\frac{1}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\log\frac{exp(z_{i}^{T}z_{i})}{exp(z_{i}^{T}z_{i})+\sum_{j\neq i}^{\mathcal{B}}exp(z_{i}^{T}z_{j})}$ (5)
@@ 3.2.2 Contrastive-based sparse compression
By leveraging the non-negative nature of latent variables $z_{i}$ in sparse autoencoders, Equation 5 can be viewed as a variant of the Non-negative Contrastive Loss (NCL) proposed in (Wang et al., 2024). This interpretation enables us to draw on the theoretical guarantees of NCL, as stated in the following theorem:
Theorem 5.
@@
(Wang et al., 2024) Under certain assumptions, the solution $\phi(x)$ is the unique solution to the NCL objective. As a result, NCL features are identifiable and disentangled.
@@
3.2.3 Training Objective
We synergistically optimize the reconstruction loss $\mathcal{L}_{\mathrm{recon}}$ and non-negative contrastive loss $\mathcal{L}_{\mathrm{ncl}}$. The former preserves semantic information with high fidelity, while the latter refines features for diverse downstream tasks. Together, these losses retain essential information and enhance the model’s discriminative power, leading to improved overall performance. The final loss function is formulated as:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{recon}}+\gamma\mathcal{L}_{\mathrm{ncl}}$ (6)

### v4

@@ Abstract
Many large-scale systems rely on high-quality deep representations (embeddings) to facilitate tasks like retrieval, search, and generative modeling. Matryoshka Representation Learning (MRL) recently emerged as a solution for adaptive embedding lengths, but it requires full model retraining and suffers from noticeable performance degradations at short lengths. In this paper, we show that sparse coding offers a compelling alternative for achieving adaptive representation with minimal overhead and higher fidelity. We propose Contrastive Sparse Representation (CSR), a method that sparsifies pre-trained embeddings into a high-dimensional but selectively activated feature space. By leveraging lightweight autoencoding and task-aware contrastive objectives, CSR preserves semantic quality while allowing flexible, cost-effective inference at different sparsity levels. Extensive experiments on image, text, and multimodal benchmarks demonstrate that CSR consistently outperforms MRL in terms of both accuracy and retrieval speed—often by large margins—while also cutting training time to a fraction of that required by MRL. Our results establish sparse coding as a powerful paradigm for adaptive representation learning in real-world applications where efficiency and fidelity are both paramount. Code is available at this https URL.
@@
Figure 1: Overview of our proposed method. (a) Illustrative comparison between standard embeddings (dense, long) and two different compression schemes: Matryoshka representations (MRL) (Kusupati et al., 2022) with short length and our Contrastive Sparse Representation (CSR) based on sparsification. (b) Comparison of retrieval accuracy and time of different methods on ImageNet with GPUs. For CSR, we present results with the SOTA RN50 backbone from Wightman (2019) as well as the same RN50 backbone from Kusupati et al. (2022) for a fair comparison. Compared to MRL and int8 quantification (Quant Int8) methods, our sparse embedding approach CSR attains the best retrieval accuracy (very close to full representations) while being much more efficient in retrieval time, using sparse matrix multiplication on GPU. (c) Training GPU hours of CSR compared to baseline methods, where we outperform MRL on average 1-NN accuracy with much less training time.
@@ 1 Introduction
Recently, a class of methods called Matryoshka Representation Learning (MRL) (Kusupati et al., 2022) has drawn a lot of attention and is now officially supported in the latest OpenAI and Google’s Gemini text embedding APIs  (OpenAI, 2024; Lee et al., 2024b) with millions of users and applications. The idea if MRL is to train an ensemble of representations truncated at different lengths (e.g., from 8 to 2048) through joint multi-task training. However, MRL deviates from standard representation learning and requires full parameter updates to the backbone; the joint training also inevitably sacrifices the quality of representations at a noticeable margin (e.g., 5% drop of top-1 accuracy on ImageNet at full representation length). These limitations render MRL a costly and lossy method for adaptive representation.
@@
Another key advantage of sparse features is that they eliminate the need to retrain the entire network. In contrast, MRL—Kusupati et al. (2022) noted—performs poorly unless full-parameter tuning. However, many existing foundation models, such as the multimodal representations in CLIP (Radford et al., 2021) and the text embeddings in NV-Embed (Lee et al., 2024a), are pre-trained as single representations on massive Internet-scale data. Fine-tuning these models would be prohibitively expensive and would prevent leveraging pre-trained open weights. Leveraging recent advances in training sparse autoencoders (SAEs) (Cunningham et al., 2023; Gao et al., 2024), we can train a lightweight 2-layer MLP module for sparsifying pre-trained embeddings within a very short period of time (e.g., half of an hour on ImageNet with a single GPU), which is of orders of magnitude faster than MRL, as shown in Figure 1(c).
These pieces of evidence on accuracy, retrieval time, and training time show that sparse features are strong alternatives to MRL methods for producing high-fidelity and computationally efficient representations with a lightweight module and training cost. Our proposed method, Contrastive Sparse Representation Learning (CSR), combines contrastive retrieval and reconstructive autoencoding objectives to preserve the original feature semantics while better tailing it down to the retrieval tasks. We evaluate CSR on a range of standard embedding benchmarks, from image embedding, text embedding, to multimodal embeddings, and compare it against various state-of-the-art efficient embedding models. Extensive experiments show that CSR consistently outperforms MRL and its variants by significant margins in terms of both accuracy and efficiency. Notably, under the same compute budget, CSR rivals MRL’s performance by 9%, 15%, and 7% on ImageNet classification, MTEB text retrieval, and MS COCO retrieval, respectively. Our main contributions are:
@@ 2 Related Work
Recent research has increasingly focused on learning adaptive representations that cater to multiple downstream tasks with diverse computational requirements. Early efforts explored context-based architectural adaptations (Kim & Cho, 2020), dynamic widths and depths in BERT (Hou et al., 2020), and random layer dropping during training to improve pruning robustness (Fan et al., 2019). More recently, Matryoshka Representation Learning (Kusupati et al., 2022) introduced a novel technique for creating flexible, nested substructures within embeddings, enabling fine-grained control over the trade-off between latency and accuracy. This concept has since been extended to various modalities and applications, including large language models (OpenAI, 2024; Nussbaum et al., 2024; Yu et al., 2024), diffusion models (Gu et al., 2023), and multimodal models (Cai et al., 2024; Hu et al., 2024). Other works have further explored token reduction in image and video processing (Yan et al., 2024b; Duggal et al., 2024).
Despite these advances, existing methods often do not fully harness the capabilities of large foundation models, highlighting the need for more effective compression strategies. Our proposed sparse compression methodology addresses this gap by providing a lightweight, plug-and-play solution that can be readily applied on top of any foundation model – significantly reducing computational overhead while preserving representational quality.
@@
Sparse coding serves as a powerful technique for compressing high-dimensional signals and extracting salient features (Wright et al., 2010; Zhang et al., 2015), with learned sparse representations often providing additional computational benefits and robustness (you2024calibrating; you2025silentmajoritydemystifyingmemorization). Prior work has induced sparsity through modifications to model design or training protocols, including modifications to attention mechanisms (Correia et al., 2019), applying Bayesian standard Gamma priors (Duan et al., 2024), incorporating discrete sparse concept layers (Koh et al., 2020; xie2025discoveringfinegrainedvisualconceptrelations), and promoting sparse activations in large language models (Mirzadeh et al., 2023; Zhang et al., 2024). However, training state-of-the-art foundation models from scratch under these sparsity constraints has proven challenging (Elhage et al., 2022), limiting their current applicability.
@@ 3 Method
Our proposed framework, Contrastive Sparse Representation learning (CSR), is illustrated in Figure 2. Starting from a pre-trained embedding $v\in\mathbb{R}^{d}$, we project it into a sparse representation space $\mathbb{R}^{h}$, selectively activating the most relevant dimensions for adaptive representation learning. We then regularize this hidden space using a reconstruction-based sparse compression loss (Section 3.2.1). Additionally, with theoretical motivations and guarantees provided by (Wang et al., 2024), we introduce a non-negative contrastive loss to expand model capacity and feature identifiability. (Section 3.2.2)
@@ 3.1 Preliminaries
Matryoshka Representation Learning (MRL).
MRL (Kusupati et al., 2022) simultaneously optimizes embeddings at multiple dimensions, as illustrated in Figure 2, to produce representations of variable size. Specifically, let $\mathcal{M}$ be a set of target embedding sizes. For each $m\in\mathcal{M}$, MRL applies an additional linear classifier to the first $m$ dimensions of the embedding vector, $v_{1:m}\in\mathbb{R}^{m}$. This design ensures each truncated representation is explicitly trained via the final loss. Formally, the MRL objective is
$\mathcal{L}_{\mathrm{MRL}}=\sum_{m\in\mathcal{M}}c_{m}\mathcal{L}_{\mathrm{CE}}\left({\bm{W}}^{(m)}\cdot f(x_{i};\theta_{f})_{1:m};y_{i}\right),$ (1)
@@
3.2 Contrastive Sparse Representation
As discussed in Section 1, MRL (Equation 1) faces two key constraints: it requires (full) training of the backbone parameters $\theta_{f}$ and its performance often deteriorates a lot under small hidden dimensions. To overcome these limitations, we propose a new methodology that relies on the computational efficiency of sparse vectors for efficient retrieval. The method, named Contrastive Sparse Representation (CSR), learns a simple one-layer sparse module on top of frozen pretrained embedding models (with full representation size, e.g., 2048) that maps dense embeddings to highly sparse embeddings with a small number of active (i.e., non-zero) dimensions (e.g., 32). As a result, CSR not only saves a lot training effort, but also allow using sparse matrix multiplication at inference time to accelerate retrieval significantly. Below, we outline how we train the CSR module through a combination of sparse autoencoding (Section 3.2.1) and sparse contrastive learning (Section 3.2.2).
3.2.1 Sparse Autoencoding
Autoencoding is a long-standing unsupervised objective that extract salient features that could preserve the original data the most a reconstruction objective. In CSR, we aim at compressing dense embeddings to sparse vectors for efficient sparse retrieval while retaining most of the useful information. To achieve this goal, we adopt sparse autoencoders due to their ability to scale with large data and restore feature semantics (Cunningham et al., 2023; Yan et al., 2024a).
Sparse Autoencoders (SAEs).
SAEs (Makhzani & Frey, 2013; Cunningham et al., 2023; Gao et al., 2024; Yan et al., 2024a) aim to extract a sparse representation $z_{k}$ by learning to reconstruct the dense feature from $z_{k}$. Specifically, given a pretrained dense embedding $v:=f(x)\in\mathbb{R}^{d}$ as the input, we apply a TopK SAE (Gao et al., 2024) with the following autoencoding process:
$z_{k}$ $:=\sigma^{+}(\operatorname{TopK}({\bm{W}}_{\mathrm{enc}}(f(x)-\bm{b}_{\mathrm{pre}})+\bm{b}_{\mathrm{enc}})),$ (2)
$\widehat{f(x)}_{k}$ $:={\bm{W}}_{\mathrm{dec}}z_{k}+\bm{b}_{\mathrm{pre}},$ (3)
where ${\bm{W}}_{\mathrm{enc}}\in\mathbb{R}^{h\times d}$ and ${\bm{W}}_{\mathrm{dec}}\in\mathbb{R}^{d\times h}$ are the encoder and decoder weight matrices, respectively; $\bm{b}_{\mathrm{enc}}\in\mathbb{R}^{h}$ and $\bm{b}_{\mathrm{pre}}\in\mathbb{R}^{d}$ are bias terms. The function $\sigma^{+}(\cdot)=\max(0,\cdot)$ denotes the ReLU activation, and $\operatorname{TopK}(\cdot)$ selects the top $k$ largest elements of the input, zeroing out the rest (as in Gao et al. (2024)). As a result, the latent $z_{k}$ is always a sparse non-negative vector with $k$ active dimensions. This enables direct control over the accuracy–compute trade-off in downstream tasks, particularly under resource-constrained conditions. We formulate the loss function as follows:
$\mathcal{L}(k)$ $=\left\|f(x)-\widehat{f(x)}_{k}\right\|_{2}^{2}.$ (4)
Moreover, as the hidden dimension $h$ increases, we empirically observe that an increasing number of latent dimensions remain inactive during training – a phenomenon referred to as “dead latents”. A large proportion of dead latents reduces the model’s capacity and leads to performance degradation (Lu et al., 2019; Templeton et al., 2024). To mitigate this issue, an auxiliary loss $\mathcal{L}_{\text{aux}}$ and Multi-TopK losses are proposed to mitigate this problem. The overall reconstruction loss is
$\mathcal{L}_{\mathrm{recon}}=\mathcal{L}(k)+\mathcal{L}(4k)/8+\beta\mathcal{L}_{\mathrm{aux}},$ (5)
@@ 3.2.1 Sparse Autoencoding
where $\mathcal{L}_{\mathrm{aux}}=||e-\hat{e}||^{2}_{2}$, $e=f(x)-\widehat{f(x)}$, and $\hat{e}=W_{\mathrm{dec}}z$ is the reconstruction using the top-$k_{\mathrm{aux}}$ dead latents. By default, we set $k_{\mathrm{aux}}=512$ and $\beta=1/32$, following the setting in Gao et al. (2024). We also offer dynamic sparsity selection, with $k$ ranging from 8 to 256, to accommodate different tasks across various modalities.
Figure 2: Overview of our proposed CSR framework. As a post-training approach, CSR differs fundamentally from MRL by projecting embeddings into a higher-dimensional space and dynamically activating only the TopK dimensions for a compact representation. The hidden space is constrained by both reconstruction and contrastive losses, which together enhance the capacity of the sparse representation while preserving computational efficiency.
3.2.2 Sparse Contrastive Learning
Furthermore, we consider to incorporate an additional sparse contrastive loss to the representations’ discriminative power. Most state-of-the-art embedding models today, e.g., CLIP (Radford et al., 2021), follow a contrastive learning paradigm, which that learns to use the embeddings to distinguish between positive and negative pairs. And it applies to both supervised and unsupervised settings (Huang et al., 2024).
The loss objective can be formulated as:
$\mathcal{L}_{\mathrm{cl}}=-\frac{1}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\log\frac{exp(z_{i}^{T}z_{i})}{exp(z_{i}^{T}z_{i})+\sum_{j\neq i}^{\mathcal{B}}exp(z_{i}^{T}z_{j})}.$ (6)
@@ 3.2.2 Sparse Contrastive Learning
By leveraging the non-negative nature of latent variables $z_{i}$ in sparse autoencoders, Equation 6 can be viewed as a variant of the Non-negative Contrastive Loss (NCL) proposed in Wang et al. (2024). This interpretation enables us to draw on the theoretical guarantees of NCL, as stated in the following theorem:
Theorem 5 (Wang et al. (2024)).
@@
Under mild conditions, the solution $\phi(x)$ is the unique solution to the NCL objective. As a result, NCL features are identifiable and disentangled.
@@
3.2.3 Overall Training Objective
At last, we optimize the sparse module through a combination of sparse autoencoding $\mathcal{L}_{\rm recon}$ and sparse contrastive learning $\mathcal{L}_{\mathrm{ncl}}$. The former incentivizes the model to preserve original semantic information in the original representation, while the latter shapes the sparse representation to be better at discriminative tasks. The final training objective of our Contrastive Sparse Representation (CSR) method is formulated as:
$\mathcal{L}_{\mathrm{CSR}}=\mathcal{L}_{\mathrm{recon}}+\gamma\mathcal{L}_{\mathrm{ncl}}.$ (7)